# Event shape dependence of symmetry plane correlations using the Gaussian estimator in Pb-Pb collisions at the LHC using AMPT

### v1

@@ III.2.1 SPCs of harmonic orders $n=2,3,6$
To make it clearer, a comparison has been shown in Fig. 4 between the SPC $\langle\cos[6(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$ which has been obtained using all events in a centrality class, to that obtained using events with higher values of $v_{6}$. We can observe no positive shift in the values, owing to the geometrical fact that the particles are not ejected with transverse momenta distributed in a hexagonal symmetry. Gauging further into the mathematical construct of the SPCs, we find that a plethora of information about the transverse momentum distribution can be derived using SPCs of planes of lower harmonic orders. Following from the discussion at the beginning of Section III, if hexagonal symmetry is dominant in the system, then $|\delta_{6}-\delta_{2}|\in\{0,\frac{\pi}{3},\frac{2\pi}{3}\}$. Thus, all possible values for $6(\delta_{6}-\delta_{2})$ are of the form $2m\pi$ for some natural number $m$. Similarly, $|\delta_{6}-\delta_{3}|\in\{0,\frac{\pi}{3}\}$, and again, $6(\delta_{6}-\delta_{3})$ will always be of the form $2m^{\prime}\pi$ for some natural number $m^{\prime}$. However, when considering a system where, let us say, a $12^{\rm th}$ order symmetry is dominant, there, $|\delta_{6}-\delta_{2}|\in\{0,\frac{\pi}{6},\frac{\pi}{3},\frac{\pi}{2},\frac{2\pi}{3}\,\frac{5\pi}{6}\}$, so that $6(\delta_{6}-\delta_{2})\in\{0,\pi,2\pi,3\pi,4\pi,5\pi\}$. So, when these pass as arguments into the cosine function in the RHS of Eq. (II.4), some terms in the numerator would be positive while some negative, so that the average in the numerator will tend towards zero, implying in turn that the symmetry planes are decorrelated. This, however, is not the case, and in order to account for the correlation between symmetry planes of harmonic orders $n=2$ and $n=3$, in a system where a $12^{\rm th}$ order symmetry dominates, we must study the SPC $\langle\cos[12(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$. If the anisotropy in the system has very high harmonic orders of symmetry planes, then that would require us to study the SPC $\lim_{k\to\infty}\langle\cos[6k(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$, which when we try to find out, starting from small natural numbers $k$, would give us vanishingly low SPC values. However, only the small $k$ values are of any physical relevance since energy occurs clustered in nuclei. It is also important to keep in mind that SPCs would invariably yield low values for an isotropic distribution of the transverse momentum vectors. In this way, a careful choice of SPCs would allow us to gauge the higher-order anisotropies of the system. It is also necessary to emphasize, however, that the errors that would be associated with finding out the dominance of higher-order symmetries in the anisotropy of the distributions are not much lesser than what we would expect by computing the anisotropic flow coefficients themselves. This is because the SPC would require a larger number of particles to be correlated as well. The usefulness of using the SPCs, however, is that we can quickly reach a conclusion about the symmetry of the dominant anisotropic flow in the system by computing each anisotropic flow coefficient and comparing them individually.
@@
Figure 5: Centrality dependence of SPCs (upper panel) for $n=2,4,6$ and corresponding PPCs (lower panel) in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT. The calculation of SPCs considers the charged hadrons with $5.00.2<p_{\rm T}<$ 5.0 GeV/c, $|\eta|<0.8$.
@@ III.2.2 SPCs of harmonic orders $n=2,4,6$
The fourth-order anisotropy would make the transverse momentum space geometry of the event appear like a square. Also, for a square, we have two distinct diagonals. These diagonals correspond to a second-order symmetry, that is, one corresponding to elliptic flow. Hence, a dominant presence of $v_{4}$ would imply that the symmetry planes $\psi_{4}$ and $\psi_{2}$ are strongly correlated. Hence, with decreasing centrality, the correlation between $\psi_{4}$ and $\psi_{2}$ is expected to increase, as is shown in Figs. 5(a) and (c). For the corresponding comparison, a plot of $v_{2}$, $v_{4}$ and $v_{6}$ is shown in Fig. 6. Furthermore, when $6^{\rm th}$ order anisotropy is dominant, it has already been described in Section III.2.1 that the correlation between $\psi_{6}$ and $\psi_{2}$ can be expected to be strong. The presence of strong hexagonal symmetry also brings with it the presence of an approximate quadrangular symmetry and, in turn, an elliptical symmetry of the anisotropy of the transverse momentum distribution of the final charged hadrons. Thus, in the presence of high $v_{4}$ and $v_{6}$, it can be expected that the symmetry planes of orders $n=2,4,6$ are well correlated, as can be seen from the comparison of Fig. 6 and Figs. 5(a) and (c).
Figure 7: Centrality dependence of SPCs (upper panel) among $n=2,3,5$ and corresponding PPCs (lower panel) in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT. The calculation of SPCs considers the charged hadrons with $5.00.2<p_{\rm T}<$ 5.0 GeV/c, $|\eta|<0.8$.
@@ III.2.3 SPCs of harmonic orders $n=2,3,5$
Whereas a pentagonal symmetry favors an approximate triangular symmetry (isosceles, but not equilateral triangles that can possibly be constructed using the vertices of a regular pentagon), the presence of elliptical symmetry is expected to be largely independent of either of the other two types of symmetries, when higher order anisotropies (e.g. $v_{30}$) are not dominant. Furthermore, as the isosceles triangles constructed out of a regular pentagon are not equilateral, we cannot expect any strong correlation between $\psi_{3}$ and $\psi_{5}$ as well. A strong SPC between harmonics $n=2,3,5$, even when different linear combinations of the symmetry plane angles are considered, would be indicative of the presence of a strong $15^{\rm th}$ order symmetry in the anisotropy of the transverse momentum distribution. As can be seen from the plots in Figs. 7(a) and (c), for two different SPCs for the same choice of harmonics, $n=2,3,5$, we see that the correlation is different, and in addition, the SPC $\langle\cos(8\psi_{2}-3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$ has values very close to zero.
@@ III.3 Spherocity dependence of SPCs and PPCs
In Figs. 35, and 7, for most of the SPCs, regardless of their centrality, it is observed that the correlators for low-$S_{0}$ (pencil-like) case have a higher magnitude of the correlation compared to $S_{0}$-int events. This is also trivially expected, because the azimuthal angles that are considered in the averaging of the multiplets in the RHS of Eq. (I) are closer to each other for low-$S_{0}$ case, and the combination $\{n_{i}\}_{i=1}^{l}$ where $\sum_{i=1}^{l}n_{i}=0$, ensures that the argument that goes in as phase into the complex exponential in the RHS of Eq. (I) is closer to zero, leading to real parts that are closer to 1. Hence, all the numerator terms during event averaging using only low-$S_{0}$ events in Eq. (II.4) will have values that are more positive in comparison to those where the event averaging is done considering all shapes of events. Thus, if an SPC computed using events of all shapes is positive in a centrality class, it would mean that the same SPC would have a higher magnitude when computed using low-$S_{0}$ events.
@@
On the other hand, when SPCs for the high-$S_{0}$ (isotropic) events, we have that the azimuthal angles are uniformly distributed, and hence the magnitude of SPCs evaluated using high-$S_{0}$ events would be smaller than that of $S_{0}$-int case. Apart from this triviality in terms of magnitude, no particular behaviour in terms of whether an SPC evaluated from isotropic events in a particular centrality class - whether it would be more positive or negative in comparison to the SPC obtained by considering events of all shapes - can be deduced from the mathematical construct of the correlators.
@@
A deviation from this trivial behaviour would be indicative of interesting underlying phenomena that lead to the geometrical distribution of particles in the transverse momentum space. The SPCs that deviate from their trivially expected values when computed using only low-$S_{0}$ or high-$S_{0}$ events are $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$, $\langle\cos(2\psi_{2}+4\psi_{4}-6\psi_{3})\rangle_{\rm GE}$, $\langle\cos[6(\psi_{2}-\psi_{3})]\rangle_{\rm GE}$ and $\langle\cos(8\psi_{2}-3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$.
@@
The participant planes are determined by the initial geometry of the colliding system. These are independent of the later stages of the evolution of the collision system. Hence, the participant planes are expected to be independent of spherocity of an event, which largely takes its value due to the later stages of evolution. Consequently, any correlation between the participant planes is also expected to be independent of the spherocity classification of an event. This is exactly what is observed in all the PPCs, which are shown in Figs. 5(d)-(f), Figs. 3(d)-(f), Figs. 7(d)-(f). The values of the PPCs match within errors for the averages that have been computed using the different kinds of event shapes. Their dependence on centrality is similar to that of the corresponding SPCs in Figs. 3(a)-(c), Figs. 5(a)-(c), Figs. 7(a)-(c).
@@
For PPCs that grow with decreasing centrality, shown in Figs. 3 (e), Fig. 5 (d) and (f), and Figs. 7 (d) and (e), a striking difference can be seen between their spherocity dependence. A noticeable feature is that the SPCs mimic qualitatively the centrality dependence of the PPCs for all correlations, at least up to mid-central events (40-50% most central). Also, there are some other SPCs, e.g, $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ and $\langle\cos(2\psi_{2}+4\psi_{4}-6\psi_{6})\rangle_{\rm GE}$ for which, the SPCs that are computed using low-$S_{0}$ events in the respective centrality class, mimic the qualitative centrality dependence of their corresponding PPCs, even for peripheral events.
@@
The general behaviour of increase of these PPCs as well as SPCs, when we move towards mid-central event classes, is that higher-order anisotropies start playing a dominant role. Hence, the lower-order symmetry planes get correlated to each other non-trivially. For most central events, that are isotropic in transverse momentum space, there need not be any correlation between the different harmonic orders of the symmetry planes, i.e., the orientations of the symmetry planes need not be correlated among themselves at all. This is exactly what is observed in both the SPCs and the PPCs, as shown in Fig. 3 (e), Fig. 5 (d) and (f), and Figs. 7 (d) and (e)
@@
Another comparison of SPCs to their respective PPCs can be made in Fig. 5(d) and Fig. 3(e), for PPCs that decrease as we move towards peripheral collisions. The reason for which $\langle\cos[6(\Phi_{3}-\Phi_{6})]\rangle_{\rm GE}$ decreases with centrality is due to its geometrical detail, as has been pointed out in Section III.2.1. For the PPC between the harmonic orders $n=2,3$ and 4, we have the fact that in most central collisions, the contribution of $v_{2}$ to anisotropic flow is less in comparison to that in peripheral events. As $\psi_{2}$ and $\psi_{4}$ together have an appreciable correlation with $\psi_{6}$ (Figs. 5(d) and (f)) and $\psi_{3}$ too has an appreciable correlation with $\psi_{6}$ (Fig. 3(d)), we observe a similar trend with centrality, for the simplest possible combination that has been constructed for the PPCs corresponding to $n=2,3$ and 4.
@@
In addition to the SPCs and PPCs discussed above, Fig. 3(f) and Fig. 7(f) show the comparison of SPCs to the corresponding PPCs which are closer to zero within error bars across all centrality classes. These correlations have very small magnitudes and vary randomly about zero. This trend applies to their respective SPCs as well. However, in these cases, it is observed that the SPCs have a higher magnitude in comparison to the PPCs, even though their variation is random about the $x-$axis. In addition, no particular behaviour in terms of event shape can be observed for the PPCs and SPCs shown in Fig. 3(c),(f) and Fig. 7(c),(f). This is evidence that the correlator is actually close to zero irrespective of the event shape chosen, and hence the SPCs obtained using averaging over pencil-like or isotropic events show no specific centrality behaviour. This behaviour of the SPCs has been reported in the ALICE experiment [23]. In all the various comparisons that were made in this section, between PPCs and their corresponding SPCs, we observe a clear qualitative match of their centrality dependence. This shows, thus, that the SPCs are a good probe for PPCs.

### v2

@@ III.2.1 SPCs of harmonic orders $n=2,3,6$
Gauging further into the mathematical construct of the SPCs, we find that a plethora of information about the transverse momentum distribution can be derived using SPCs of planes of lower harmonic orders. Following from the discussion at the beginning of Section III, if hexagonal symmetry is dominant in the system, then $|\delta_{6}-\delta_{2}|\in\{0,\frac{\pi}{3},\frac{2\pi}{3}\}$. Thus, all possible values for $6(\delta_{6}-\delta_{2})$ are of the form $2m\pi$ for some natural number $m$. Similarly, $|\delta_{6}-\delta_{3}|\in\{0,\frac{\pi}{3}\}$, and again, $6(\delta_{6}-\delta_{3})$ will always be of the form $2m^{\prime}\pi$ for some natural number $m^{\prime}$. However, when considering a system where, let us say, a $12^{\rm th}$ order symmetry is dominant, there, $|\delta_{6}-\delta_{2}|\in\{0,\frac{\pi}{6},\frac{\pi}{3},\frac{\pi}{2},\frac{2\pi}{3},\frac{5\pi}{6}\}$, so that $6(\delta_{6}-\delta_{2})\in\{0,\pi,2\pi,3\pi,4\pi,5\pi\}$. So, when these pass as arguments into the cosine function in the RHS of Eq. (5), some terms in the numerator would be positive while some negative, so that the average in the numerator will tend towards zero, implying in turn that the symmetry planes are decorrelated. This, however, is not the case, and in order to account for the correlation between symmetry planes of harmonic orders $n=2$ and $n=3$, in a system where a $12^{\rm th}$ order symmetry dominates, we must study the SPC $\langle\cos[12(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$.
In general, if the anisotropy in the system has very high harmonic orders of symmetry planes, then that would require us to study the SPC $\lim_{k\to\infty}\langle\cos[6k(\psi_{3}-\psi_{2})]\rangle_{\rm GE}$, which when we try to find out, starting from small natural numbers $k$, would give us vanishingly low SPC values. However, only the small $k$ values are of any physical relevance since energy occurs clustered in nuclei. It is also important to keep in mind that SPCs would invariably yield low values for an isotropic distribution of the transverse momentum vectors. In this way, a careful choice of SPCs would allow us to gauge the higher-order anisotropies of the system. It is also necessary to emphasize, however, that the errors that would be associated with finding out the dominance of higher-order symmetries in the anisotropy of the distributions are not much lesser than what we would expect by computing the anisotropic flow coefficients themselves. This is because the SPC would require a larger number of particles to be correlated as well. The usefulness of using the SPCs, however, is that we can quickly reach a conclusion about the symmetry of the dominant anisotropic flow in the system, rather than by computing each anisotropic flow coefficient and comparing them individually.
@@
Figure 5: Centrality dependence of SPCs (shown in (a), (b), and (c)) for $n=2,4,6$ and corresponding PPCs (shown in (d), (e), and (f)) in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT. The calculation of SPCs considers the charged hadrons with $0.2<p_{\rm T}<5.0$ GeV/c, $|\eta|<0.8$.
@@ III.2.2 SPCs of harmonic orders $n=2,4,6$
The fourth-order anisotropy would make the transverse momentum space geometry of the event appear like a square. Also, for a square, we have two distinct diagonals. These diagonals correspond to a second-order symmetry, that is, one corresponding to elliptic flow. Hence, a dominant presence of $v_{4}$ would imply that the symmetry planes $\psi_{4}$ and $\psi_{2}$ are strongly correlated. Hence, with decreasing centrality, the correlation between $\psi_{4}$ and $\psi_{2}$ is expected to increase, as has also been seen in Figs. 5(a) and (c). For the corresponding comparison, a plot of $v_{2}$, $v_{4}$ and $v_{6}$ is shown in Fig. 6. Furthermore, when $6^{\rm th}$ order anisotropy is dominant, it has already been described in Section III.2.1 that the correlation between $\psi_{6}$ and $\psi_{2}$ can be expected to be strong. The presence of strong hexagonal symmetry also brings with it the presence of an approximate quadrangular symmetry and, in turn, an elliptical symmetry of the anisotropy of the transverse momentum distribution of the final charged hadrons. Thus, in the presence of high $v_{4}$ and $v_{6}$, it can be expected that the symmetry planes of orders $n=2,4,6$ are well correlated, as can be seen from the comparison of Fig. 6 and Figs. 5(a) and (c).
Figure 7: Centrality dependence of SPCs (shown in (a), (b), and (c)) among $n=2,3,5$ and corresponding PPCs (shown in (d), (e), and (f)) in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV using AMPT. The calculation of SPCs considers the charged hadrons with $0.2<p_{\rm T}<5.0$ GeV/c, $|\eta|<0.8$.
@@ III.2.3 SPCs of harmonic orders $n=2,3,5$
Whereas a pentagonal symmetry favors an approximate triangular symmetry (not equilateral, but isosceles triangles that can possibly be constructed using the vertices of a regular pentagon), the presence of elliptical symmetry is expected to be largely independent of either of the other two symmetries, when higher order anisotropies (e.g. $v_{30}$) are not dominant. However, the SPCs are also expected to arise from the cumulant expansion of the harmonics, where the higher-order harmonics have nonlinear contributions from the lower-order harmonics [50, 65, 66, 58]. This explains the positive correlation the symmetry planes $\psi_{a}$, $\psi_{b}$, and $\psi_{c}$ for $c>b>a$ and $na+mb=c$ with $n,~m\in\mathbb{N}$. This is explicitly observed in the strong positive correlation in $\langle\cos(2\psi_{2}+3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$ (Fig. 7 (a)) which is absent in $\langle\cos[2\psi_{2}+4\psi_{4}-6\psi_{3}]\rangle_{\rm GE}$ (Fig. 5 (b)). It is to be noted that a weaker correlation for $\langle\cos[8\psi_{2}-3\psi_{3}-5\psi_{5}]\rangle_{\rm GE}$ than $\langle\cos[2\psi_{2}+3\psi_{3}-5\psi_{5}]\rangle_{\rm GE}$ is attributed to the higher number of particle correlations in the former case.
For most of the correlators, it can be observed that the transformation from PPCs to SPCs is non-linear in nature. The non-linear response can be attributed to the following factors.
  1. 1.
    As one moves from central to peripheral collisions, the number of participants (and parton density) decreases, which makes it difficult to transform initial PPCs to SPCs.
  2. 2.
    Towards the peripheral collisions, due to a smaller number of participants and charged particle multiplicities, the event-by-event fluctuations dominate the values of PPCs. A non-linear contribution of these lower order PPCs to the higher order SPCs can induce a overall non-linear response of initial PPCs to final state SPCs.
  3. 3.
    The Gaussian estimator, although better than other estimators of SPCs, is still not robust to a change in multiplicity.
A similar non-linear response of the system to the evolution of participant eccentricities to the final state anisotropic flow coefficients is also shown in Refs. [59, 33].
@@ III.3 Spherocity dependence of SPCs and PPCs
In Figs. 35, and 7, for most of the SPCs, regardless of their centrality, it is observed that the correlators for low-$S_{0}$ events have a higher magnitude of the correlation compared to $S_{0}$-int events. This is also true for the cases where the SPCs are driven by the elliptic geometry directly or indirectly arising in the cumulant expansion of the higher-order harmonics. This is because, in low-$S_{0}$ events, the azimuthal angles of a larger number of emergent particles are aligned to a particular direction in momentum space, which is characterised by the parametric direction $\hat{n}$ in Eq. 4, than in the case of an event with high-$S_{0}$ of similar multiplicity. Therefore, in comparison to other events, since the azimuthal angles of the emergent particles are closer together in the momentum space for a low $S_{0}$ event, the combination $\{n_{i}\}_{i=1}^{l}$ where $\sum_{i=1}^{l}n_{i}=0$, ensures that the argument that goes in as phase into the complex exponential in the RHS of Eq. (I) is closer to zero. This implies that the low-$S_{0}$ events have dominating elliptic symmetry, and higher order harmonics, including triangular flow, contribute negligibly. The elliptic symmetry in the azimuthal distribution of particles ensures a vanishing sine term and a large cosine term in the RHS of Eq. (I). This leads to larger positive real parts. Hence, all the numerator terms during event averaging using only low-$S_{0}$ events in Eq. (5) will have values that are more positive in comparison to those where the event averaging is done considering all shapes of events. Thus, if an SPC computed using events of all shapes is positive in a centrality class, it would mean that the same SPC would have a higher magnitude when computed using low-$S_{0}$ events.
@@
On the other hand, when SPCs for the high-$S_{0}$ (isotropic) events are considered, we have that the azimuthal angles are uniformly distributed, and hence the magnitude of SPCs evaluated using high-$S_{0}$ events would be smaller than that of the $S_{0}$-int case. Apart from this triviality in terms of magnitude, no particular behaviour in terms of whether an SPC evaluated from isotropic events in a particular centrality class - if it would be more positive or negative in comparison to the SPC obtained by considering events of all shapes - can be deduced solely from the mathematical construct of the correlators.
@@
A deviation from this trivial behaviour would be indicative of interesting underlying phenomena that lead to the observed geometrical distribution of particles in the transverse momentum space. The SPCs that deviate from their trivially expected values when computed using only low-$S_{0}$ or high-$S_{0}$ events are $\langle\cos[6(\psi_{6}-\psi_{3})]\rangle_{\rm GE}$, $\langle\cos(2\psi_{2}+4\psi_{4}-6\psi_{3})\rangle_{\rm GE}$, $\langle\cos[6(\psi_{2}-\psi_{3})]\rangle_{\rm GE}$ and $\langle\cos(8\psi_{2}-3\psi_{3}-5\psi_{5})\rangle_{\rm GE}$.
@@
The participant planes are determined by the initial geometry of the colliding system. These are independent of the later stages of the evolution of the collision system. Hence, the participant planes are expected to be independent of the spherocity of an event, which largely takes its value due to the later stages of evolution. Consequently, any correlation between the participant planes is also expected to be independent of the spherocity classification of an event. This is exactly what is observed in all the PPCs, which are shown in Figs. 5(d)-(f), Figs. 3(d)-(f), Figs. 7(d)-(f). The values of the PPCs match within errors for the averages that have been computed using the different kinds of event shapes. Their dependence on centrality is similar to that of the corresponding SPCs in Figs. 3(a)-(c), Figs. 5(a)-(c), Figs. 7(a)-(c).
@@
For PPCs that grow with decreasing centrality, shown in Figs. 3 (e), Fig. 5 (d) and (f), and Figs. 7 (d) and (e), a striking difference can be seen between their spherocity dependence. A noticeable feature is that the SPCs mimic qualitatively the centrality dependence of the PPCs for all correlations, at least up to mid-central events (40-50% most central). Also, there are some other SPCs, e.g, $\langle\cos[6(\psi_{6}-\psi_{2})]\rangle_{\rm GE}$ and $\langle\cos(2\psi_{2}+4\psi_{4}-6\psi_{6})\rangle_{\rm GE}$ for which the SPCs that are computed using low-$S_{0}$ events in the respective centrality class, mimic the qualitative centrality dependence of their corresponding PPCs, even for peripheral events.
@@
The general behaviour of the increase of these PPCs as well as SPCs, when we move towards mid-central event classes, is that higher-order anisotropies start playing a dominant role. Hence, the lower-order symmetry planes get correlated to each other non-trivially. For most central events, which are isotropic in transverse momentum space, there need not be any correlation between the different harmonic orders of the symmetry planes, i.e., the orientations of the symmetry planes need not be correlated among themselves at all. This is exactly what is observed in both the SPCs and the PPCs, as shown in Fig. 3 (e), Fig. 5 (d) and (f), and Figs. 7 (d) and (e).
@@
Another comparison of SPCs to their respective PPCs can be made in Fig. 5(e) and Fig. 3(d), for PPCs that decrease as we move towards peripheral collisions. The reason for which $\langle\cos[6(\Phi_{3}-\Phi_{6})]\rangle_{\rm GE}$ decreases with centrality is due to its geometrical detail, as has been pointed out in Section III.2.1. For the PPC between the harmonic orders $n=2,3$ and 4, we have the fact that in most central collisions, the contribution of $v_{2}$ to anisotropic flow is less in comparison to that in peripheral events. As $\psi_{2}$ and $\psi_{4}$ together have an appreciable correlation with $\psi_{6}$ (Figs. 5(d) and (f)) and $\psi_{3}$ too has an appreciable correlation with $\psi_{6}$ (Fig. 3(d)), we observe a similar trend with centrality, for the simplest possible combination that has been constructed for the PPCs corresponding to $n=2,3$ and 4.
@@
In addition to the SPCs and PPCs discussed above, Fig. 3(f) and Fig. 7(f) show the comparison of SPCs to the corresponding PPCs, which are closer to zero within error bars across all centrality classes. These correlations have very small magnitudes and vary randomly about zero. This trend applies to their respective SPCs as well. However, in these cases, it is observed that the SPCs have a higher magnitude in comparison to the PPCs, even though their variation is random about the $x$-axis. In addition, no particular behaviour in terms of event shape can be observed for the PPCs and SPCs shown in Fig. 3(c),(f) and Fig. 7(c),(f). This is an indication that the correlator is actually close to zero, irrespective of the event shape chosen, and hence the SPCs obtained using averaging over pencil-like or isotropic events show no specific centrality behaviour. This behaviour of the SPCs has been reported in the ALICE experiment [23]. In all the various comparisons that were made in this section, between PPCs and their corresponding SPCs, we observe a clear qualitative match of their centrality dependence. This shows, thus, that the SPCs are a good probe for PPCs.